# Simple Policy Optimization

### v7

@@ Abstract
Figure 1: Training performance in the Breakout environment. SPO is the first model-free algorithm capable of end-to-end training for extremely deep neural network architectures, positioning itself as a promising alternative to the well-known PPO algorithm.
@@ 1 Introduction
Deep Reinforcement Learning (DRL) has achieved great success in recent years, notably in board games (Silver et al., 2016, 2017, 2018), video games (Mnih et al., 2015; Berner et al., 2019; Vinyals et al., 2019; Ye et al., 2020), autonomous driving (Kiran et al., 2021; Zhu & Zhao, 2021; Teng et al., 2023), and robotics (Makoviychuk et al., 2021; Brunke et al., 2022; Han et al., 2023). Policy gradient (PG) methods (Sutton & Barto, 2018; Lehmann, 2024), as a major paradigm in RL, have been widely adopted by the academic community. One main practical challenge of PG methods is to reduce the variance of the gradients while keeping the bias low (Sutton et al., 2000; Schulman et al., 2015b). In this context, a widely used technique is to add a baseline (usually the value function) when sampling an estimate of the action-value function (Greensmith et al., 2004). Another challenge of PG methods is to estimate the proper step size for the policy update (Kakade & Langford, 2002; Schulman et al., 2015a). Given that the training data strongly depends on the current policy, a large step size may result in a collapse of policy performance, whereas a small one may impair the sample efficiency of the algorithm.
@@
To address these challenges, in Trust Region Policy Optimization (TRPO), Schulman et al. (2015a) proved that minimizing a certain surrogate objective guarantees policy improvement with non-trivial step sizes. Subsequently, the TRPO algorithm was derived through a series of approximations, which impose a trust region constraint during the policy iterations, leading to monotonic policy improvement in theory. However, given the complexity of second-order optimization, TRPO is highly inefficient and can be hard to extend to large-scale RL environments. Proximal Policy Optimization (PPO) (Schulman et al., 2017) is designed to enforce comparable constraints on the difference between successive policies during the training process, while only using first-order optimization. By clipping the current data that exceeds the probability ratio limit to a constant, PPO attempts to remove the high incentive for pushing the current policy away from the old one. It has been shown that PPO can be effectively extended to large-scale complex control tasks (Ye et al., 2020; Makoviychuk et al., 2021).
Despite its success, the optimization behavior of PPO remains insufficiently understood. Although PPO aims to constrain the probability ratio deviations between successive policies, it often fails to keep these ratios within bounds (Ilyas et al., 2018; Engstrom et al., 2020). In some tasks, the ratios can even escalate to values as high as $40$ (Wang et al., 2020). Furthermore, studies have revealed that PPO’s performance is highly dependent on “code-level optimizations” (Andrychowicz et al., 2021). The implementation of PPO includes numerous code-level details that critically influence its effectiveness (Engstrom et al., 2020; Huang et al., 2022a, b).
In this paper, we propose a new algorithm named Simple Policy Optimization (SPO) designed to more effectively bound probability ratios through a novel objective function. The differences in optimization behavior between PPO and SPO are illustrated in Figure 2. Our main contributions are summarized as follows:
@@ 2 Related Work
However, empirical evidence from a wide range of studies demonstrates that ratio clipping fails to effectively enforce trust region constraints (Wang et al., 2020). To prevent aggressive policy updates, some works have focused on designing adaptive learning rates based on TV divergence or KL divergence (Heess et al., 2017; Queeney et al., 2021; Rudin et al., 2022), which have been shown to effectively enhance the stability of PPO. On the other hand, code-level optimizations are crucial for the robust performance of PPO (Engstrom et al., 2020). High-quality implementations of PPO involve numerous code details (Huang et al., 2022a, b), making it challenging to accurately assess the core factors that truly affect the algorithm’s performance.
In this work, we show that heuristic clipping technique can not enforce trust region constraints (see Figure 2). During PPO’s iterations, ratio clipping zeros the gradients of certain data points, which can lead to a lack of corrective gradients to prevent the policy from escaping the trust region, thus undermining the monotonic improvement guarantee. As a result, PPO requires additional code-level tuning, such as adaptive learning rates or early stopping strategies, to artificially prevent performance collapse. We reveal this inherent flaw of ratio clipping and propose promising alternatives.
@@ 3.2 Trust Region Policy Optimization
$\begin{split}\eta(\tilde{\pi})-\eta(\pi)\geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)},a\sim{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\pi}(\cdot|s)}}\left[A_{\pi}(s,a)\right]\\ &-\frac{2\epsilon\gamma}{(1-\gamma)^{2}}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)}}\left[D_{\mathrm{TV}}(\pi\|\tilde{\pi})[s]\right],\\ \end{split}$ (3)
where $\epsilon=\max_{s}\left|\mathbb{E}_{a\sim\tilde{\pi}(\cdot|s)}\left[A_{\pi}(s,a)\right]\right|$, $D_{\mathrm{TV}}$ is the Total Variation (TV) divergence.
@@
$\begin{split}\eta(\tilde{\pi})-\eta(\pi)\geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)},a\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\pi(\cdot|s)}}\left[\frac{\tilde{\pi}(a|s)}{\pi(a|s)}\cdot A_{\pi}(s,a)\right]\\ &-\frac{2\epsilon\gamma}{(1-\gamma)^{2}}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)}}\left[\sqrt{\frac{1}{2}D_{\mathrm{KL}}(\pi\|\tilde{\pi})[s]}\right].\\ \end{split}$ (4)
@@ 4 Methodology
$\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)}}\left[D_{\mathrm{TV}}(\pi\|\tilde{\pi})[s]\right]=\frac{1}{2}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)},a\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\pi(\cdot|s)}}\left|\frac{\tilde{\pi}(a|s)}{\pi(a|s)}-1\right|.$ (9)
@@
$\begin{split}\eta(\tilde{\pi})-\eta(\pi)\geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)},a\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\pi(\cdot|s)}}\left[\frac{\tilde{\pi}(a|s)}{\pi(a|s)}\cdot A_{\pi}(s,a)\right]\\ &-\frac{\epsilon\gamma}{(1-\gamma)^{2}}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)},a\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\pi(\cdot|s)}}\left|\frac{\tilde{\pi}(a|s)}{\pi(a|s)}-1\right|.\\ \end{split}$ (10)
@@
Given any state $s\in\mathcal{S}$ and the old policy $\pi$, define the solution spaces under the TV and KL divergence constraints as follows:
$\begin{split}\Omega_{\mathrm{TV}}&=\{\tilde{\pi}\mid D_{\mathrm{TV}}\left[\pi\|\tilde{\pi}\right][s]\leq\delta_{\mathrm{TV}},\forall s\in\mathcal{S}\},\\ \Omega_{\mathrm{KL}}&=\{\tilde{\pi}\mid D_{\mathrm{KL}}\left[\pi\|\tilde{\pi}\right][s]\leq\delta_{\mathrm{KL}},\forall s\in\mathcal{S}\},\\ \end{split}$ (11)
where $\delta_{\mathrm{KL}}>0$ is a predefined threshold. Under these constraints, there exists a map $f$ such that, for any given $\delta_{\mathrm{KL}}$, let $\delta_{\mathrm{TV}}=f(\delta_{\mathrm{KL}})$, we establish that $\Omega_{\mathrm{KL}}\subset\Omega_{\mathrm{TV}}$.
@@
Let $\delta_{\mathrm{TV}}=\sqrt{\frac{1}{2}\delta_{\mathrm{KL}}}$. For any $s\in\mathcal{S}$ and $\tilde{\pi}\in\Omega_{\mathrm{KL}}$, using Pinsker’s inequality, we have $D_{\mathrm{TV}}\left[\pi\|\tilde{\pi}\right][s]\leq\sqrt{\frac{1}{2}D_{\mathrm{KL}}\left[\pi\|\tilde{\pi}\right][s]}\leq\sqrt{\frac{1}{2}\delta_{\mathrm{KL}}}=\delta_{\mathrm{TV}}$, therefore $\tilde{\pi}\in\Omega_{\mathrm{TV}}$, which means $\Omega_{\mathrm{KL}}\subset\Omega_{\mathrm{TV}}$, concluding the proof. ∎
@@
$\begin{split}\mathcal{L}_{\pi}^{\mathrm{TV}}(\tilde{\pi})=&\frac{1}{1-\gamma}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)},a\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\pi(\cdot|s)}}\left[\frac{\tilde{\pi}(a|s)}{\pi(a|s)}\cdot A_{\pi}(s,a)\right]\\ &-\frac{2\epsilon\gamma}{(1-\gamma)^{2}}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)}}\left[D_{\mathrm{TV}}(\pi\|\tilde{\pi})[s]\right],\\ \mathcal{L}_{\pi}^{\mathrm{KL}}(\tilde{\pi})=&\frac{1}{1-\gamma}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)},a\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\pi(\cdot|s)}}\left[\frac{\tilde{\pi}(a|s)}{\pi(a|s)}\cdot A_{\pi}(s,a)\right]\\ &-\frac{2\epsilon\gamma}{(1-\gamma)^{2}}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)}}\left[\sqrt{\frac{1}{2}D_{\mathrm{KL}}(\pi\|\tilde{\pi})[s]}\right].\\ \end{split}$ (12)
@@
$\begin{split}\mathcal{L}_{\pi}^{\mathrm{TV}}&(\tilde{\pi}_{\mathrm{TV}}^{*})\geq\mathcal{L}_{\pi}^{\mathrm{TV}}(\tilde{\pi}_{\mathrm{KL}}^{*})=\\ &\frac{1}{1-\gamma}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)},a\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\pi(\cdot|s)}}\left[\frac{\tilde{\pi}_{\mathrm{KL}}^{*}(a|s)}{\pi(a|s)}\cdot A_{\pi}(s,a)\right]\\ &-\frac{2\epsilon\gamma}{(1-\gamma)^{2}}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)}}\left[D_{\mathrm{TV}}(\pi\|\tilde{\pi}_{\mathrm{KL}}^{*})[s]\right]\geq\\ &\frac{1}{1-\gamma}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)},a\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\pi(\cdot|s)}}\left[\frac{\tilde{\pi}_{\mathrm{KL}}^{*}(a|s)}{\pi(a|s)}\cdot A_{\pi}(s,a)\right]\\ &-\frac{2\epsilon\gamma}{(1-\gamma)^{2}}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)}}\left[\sqrt{\frac{1}{2}D_{\mathrm{KL}}(\pi\|\tilde{\pi}_{\mathrm{KL}}^{*})[s]}\right]\\ =&\mathcal{L}_{\pi}^{\mathrm{KL}}(\tilde{\pi}_{\mathrm{KL}}^{*}),\\ \end{split}$ (14)
concluding the proof. ∎
@@
As a result, we aim to solve the following constrained optimization problem:
@@
PPO attempts to satisfy the constraints of (15) through ratio clipping, but this does not prevent excessive ratio deviations (see Figure 2). The underlying reason is that ratio clipping causes certain data points to stop contributing to the gradients. Over multiple iterations, this can lead to uncontrollable updates, as the absence of corrective gradients prevents the policy from recovering. To overcome this issue with ratio clipping, we propose the following objective:

### v8

@@ Abstract
Figure 1: Training performance in the Breakout environment. SPO is a novel model-free algorithm capable of end-to-end training for extremely deep neural network architectures, positioning itself as a promising alternative to the well-known PPO algorithm.
@@ 1 Introduction
Deep Reinforcement Learning (DRL) has achieved great success in recent years, notably in games (Mnih et al., 2015; Silver et al., 2016, 2017, 2018; Vinyals et al., 2019), foundation model fine-tuning (Ouyang et al., 2022; Black et al., 2023), and robotic control (Makoviychuk et al., 2021; Rudin et al., 2022). Policy gradient (PG) methods (Sutton & Barto, 2018; Lehmann, 2024), as a major paradigm in RL, have been widely adopted by the academic community. One main practical challenge of PG methods is to reduce the variance of the gradients while keeping the bias low (Sutton et al., 2000; Schulman et al., 2015b). In this context, a widely used technique is to add a baseline when sampling an estimate of the action-value function (Greensmith et al., 2004). Another challenge of PG methods is to estimate the proper step size for the policy update (Kakade & Langford, 2002; Schulman et al., 2015a). Given that the training data strongly depends on the current policy, a large step size may result in a collapse of policy performance, whereas a small one may impair the sample efficiency of the algorithm.
@@
To address these challenges, Schulman et al. (2015a) proved that optimizing a certain surrogate objective guarantees policy improvement with non-trivial step sizes. Subsequently, the TRPO algorithm was derived through a series of approximations, which impose a trust region constraint during the policy iterations, leading to monotonic policy improvement in theory. However, given the complexity of second-order optimization, TRPO is highly inefficient and can be hard to extend to large-scale RL environments. Proximal Policy Optimization (PPO) (Schulman et al., 2017) is designed to enforce comparable constraints on the difference between successive policies during the training process, while only using first-order optimization. By clipping the current data that exceeds the probability ratio limit to a constant, PPO attempts to remove the high incentive for pushing the current policy away from the old one. It has been demonstrated that PPO can be effectively extended to large-scale complex control tasks (Ye et al., 2020; Makoviychuk et al., 2021).
Despite its success, the optimization behavior of PPO remains insufficiently understood. Although PPO aims to constrain the probability ratio deviations between successive policies, it often fails to keep these ratios within bounds (Ilyas et al., 2018; Engstrom et al., 2020; Wang et al., 2020). In some tasks, the ratios can even escalate to values as high as $40$ (Wang et al., 2020). Furthermore, studies have revealed that PPO’s performance is highly dependent on “code-level optimizations” (Andrychowicz et al., 2021; Huang et al., 2022a). The implementation of PPO includes numerous code-level details that critically influence its effectiveness (Engstrom et al., 2020; Huang et al., 2022b).
In this paper, we propose a new model-free RL algorithm named Simple Policy Optimization (SPO) designed to more effectively bound probability ratios through a novel objective function. The key differences in optimization behavior between PPO and SPO are illustrated in Figure 2. Our main contributions are summarized as follows:
@@ 2 Related Work
However, empirical evidence from a wide range of studies demonstrates that ratio clipping fails to enforce trust region constraints effectively (Wang et al., 2020). To prevent aggressive policy updates, previous works have focused on designing adaptive learning rates based on TV divergence or KL divergence (Heess et al., 2017; Queeney et al., 2021; Rudin et al., 2022), which have been shown to effectively enhance the stability of PPO. On the other hand, code-level optimizations are crucial for the robust performance of PPO (Engstrom et al., 2020). High-quality implementations of PPO involve numerous code details (Huang et al., 2022a, b), making it challenging to accurately assess the core factors that truly affect the algorithm’s performance.
In this work, we argue that heuristic clipping technique cannot enforce trust region constraints (see Figure 2). During PPO’s iterations, ratio clipping zeros the gradients of certain data points, which can lead to a lack of corrective gradients to prevent the policy from escaping the trust region, thus undermining the monotonic improvement guarantee. As a result, PPO requires additional code-level tuning, such as adaptive learning rates or early stopping strategies, to artificially prevent performance collapse. We reveal this inherent flaw of ratio clipping and propose promising alternatives.
@@ 3.2 Trust Region Policy Optimization
$\begin{split}\eta(\tilde{\pi})-\eta(\pi)\geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)},a\sim{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\pi}(\cdot|s)}}\left[A_{\pi}(s,a)\right]\\ &-\frac{2\xi\gamma}{(1-\gamma)^{2}}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)}}\left[D_{\mathrm{TV}}(\pi\|\tilde{\pi})[s]\right],\\ \end{split}$ (3)
where $\xi=\max_{s}\left|\mathbb{E}_{a\sim\tilde{\pi}(\cdot|s)}\left[A_{\pi}(s,a)\right]\right|$, $D_{\mathrm{TV}}$ is the Total Variation (TV) divergence.
@@
$\begin{split}\eta(\tilde{\pi})-\eta(\pi)\geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)},a\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\pi(\cdot|s)}}\left[\frac{\tilde{\pi}(a|s)}{\pi(a|s)}\cdot A_{\pi}(s,a)\right]\\ &-\frac{2\xi\gamma}{(1-\gamma)^{2}}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)}}\left[\sqrt{\frac{1}{2}D_{\mathrm{KL}}(\pi\|\tilde{\pi})[s]}\right].\\ \end{split}$ (4)
@@ 4 Methodology
$\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)}}\left[D_{\mathrm{TV}}(\pi\|\tilde{\pi})[s]\right]=\frac{1}{2}\mathop{\mathbb{E}}_{\begin{subarray}{c}s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)}\\ a\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\pi(\cdot|s)}\end{subarray}}\left[\left|\frac{\tilde{\pi}(a|s)}{\pi(a|s)}-1\right|\right].$ (9)
@@
$\begin{split}\eta(\tilde{\pi})-\eta(\pi)\geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)},a\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\pi(\cdot|s)}}\left[\frac{\tilde{\pi}(a|s)}{\pi(a|s)}\cdot A_{\pi}(s,a)\right]\\ -&\frac{\xi\gamma}{(1-\gamma)^{2}}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)},a\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\pi(\cdot|s)}}\left[\left|\frac{\tilde{\pi}(a|s)}{\pi(a|s)}-1\right|\right].\\ \end{split}$ (10)
@@
Given the old policy $\pi$, define the solution spaces under the TV and KL divergence constraints as follows:
$\begin{split}\Omega_{\mathrm{TV}}&=\{\tilde{\pi}\mid D_{\mathrm{TV}}(\pi\|\tilde{\pi})[s]\leq\delta_{\mathrm{TV}},\forall s\in\mathcal{S}\},\\ \Omega_{\mathrm{KL}}&=\{\tilde{\pi}\mid D_{\mathrm{KL}}(\pi\|\tilde{\pi})[s]\leq\delta_{\mathrm{KL}},\forall s\in\mathcal{S}\},\\ \end{split}$ (11)
where $\delta_{\mathrm{KL}}>0$ is a predefined threshold. Let $\delta_{\mathrm{TV}}=\sqrt{\frac{1}{2}\delta_{\mathrm{KL}}}$, we establish that $\Omega_{\mathrm{KL}}\subset\Omega_{\mathrm{TV}}$.
@@
For any given $\delta_{\mathrm{KL}}$ and $\tilde{\pi}\in\Omega_{\mathrm{KL}}$, using Pinsker’s inequality, we have $D_{\mathrm{TV}}(\pi\|\tilde{\pi})[s]\leq\sqrt{\frac{1}{2}D_{\mathrm{KL}}(\pi\|\tilde{\pi})[s]}\leq\sqrt{\frac{1}{2}\delta_{\mathrm{KL}}}=\delta_{\mathrm{TV}}$, therefore $\tilde{\pi}\in\Omega_{\mathrm{KL}}\Longrightarrow\tilde{\pi}\in\Omega_{\mathrm{TV}}$, which means $\Omega_{\mathrm{KL}}\subset\Omega_{\mathrm{TV}}$, concluding the proof. ∎
@@
$\begin{split}\mathcal{L}_{\pi}^{\mathrm{TV}}(\tilde{\pi})=&\frac{1}{1-\gamma}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)},a\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\pi(\cdot|s)}}\left[\frac{\tilde{\pi}(a|s)}{\pi(a|s)}\cdot A_{\pi}(s,a)\right]\\ &-\frac{2\xi\gamma}{(1-\gamma)^{2}}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)}}\left[D_{\mathrm{TV}}(\pi\|\tilde{\pi})[s]\right],\\ \mathcal{L}_{\pi}^{\mathrm{KL}}(\tilde{\pi})=&\frac{1}{1-\gamma}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)},a\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\pi(\cdot|s)}}\left[\frac{\tilde{\pi}(a|s)}{\pi(a|s)}\cdot A_{\pi}(s,a)\right]\\ &-\frac{2\xi\gamma}{(1-\gamma)^{2}}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)}}\left[\sqrt{\frac{1}{2}D_{\mathrm{KL}}(\pi\|\tilde{\pi})[s]}\right].\\ \end{split}$ (12)
@@
$\begin{split}\mathcal{L}_{\pi}^{\mathrm{TV}}&(\tilde{\pi}_{\mathrm{TV}}^{*})\geq\mathcal{L}_{\pi}^{\mathrm{TV}}(\tilde{\pi}_{\mathrm{KL}}^{*})=\\ &\frac{1}{1-\gamma}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)},a\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\pi(\cdot|s)}}\left[\frac{\tilde{\pi}_{\mathrm{KL}}^{*}(a|s)}{\pi(a|s)}\cdot A_{\pi}(s,a)\right]\\ &-\frac{2\xi\gamma}{(1-\gamma)^{2}}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)}}\left[D_{\mathrm{TV}}(\pi\|\tilde{\pi}_{\mathrm{KL}}^{*})[s]\right]\geq\\ &\frac{1}{1-\gamma}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)},a\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\pi(\cdot|s)}}\left[\frac{\tilde{\pi}_{\mathrm{KL}}^{*}(a|s)}{\pi(a|s)}\cdot A_{\pi}(s,a)\right]\\ &-\frac{2\xi\gamma}{(1-\gamma)^{2}}\mathbb{E}_{s\sim{\color[rgb]{0,120,255}\definecolor[named]{pgfstrokecolor}{rgb}{0,120,255}\rho_{\pi}(\cdot)}}\left[\sqrt{\frac{1}{2}D_{\mathrm{KL}}(\pi\|\tilde{\pi}_{\mathrm{KL}}^{*})[s]}\right]\\ =&\mathcal{L}_{\pi}^{\mathrm{KL}}(\tilde{\pi}_{\mathrm{KL}}^{*}),\\ \end{split}$ (14)
thus $\mathcal{L}_{\pi}^{\mathrm{TV}}(\tilde{\pi}_{\mathrm{TV}}^{*})\geq\mathcal{L}_{\pi}^{\mathrm{KL}}(\tilde{\pi}_{\mathrm{KL}}^{*})$, concluding the proof. ∎
@@
As a result, to optimize the lower bound (10), we aim to solve the following constrained optimization problem:
@@
PPO attempts to satisfy the constraints of (15) through ratio clipping, but this does not prevent excessive ratio deviations (demonstrated in Figure 2). The underlying reason is that ratio clipping causes certain data points to stop contributing to the gradients. Over multiple iterations, this can lead to uncontrollable updates, as the absence of corrective gradients prevents the policy from recovering. To overcome this issue with ratio clipping, we propose the following objective: